Your friend is taking an algebra class that began one week after ours. Explain to your friend how by looking at two linear equations you can tell that corresponding lines are parallel and how you can tell that corresponding lines are perpendicular from the equations alone.
Answer by mananth(16946)

write the equations in the form
y=mx +c
m is the slope of the line.
Let the slope of one line be m1
Then the equation will be
y=m1x+c
..
Let the slope of the second line be m2
Write the equation of the second line in the same form
y=m2x+c
..
If m1 * m2 = -1 then the lines are perpendicular
OR m1 = -1/m2